Consider the following System of equations:

3π‘₯ + 2𝑦 + 𝑧 = 3

2π‘₯ + 𝑦 + 𝑧 = 0

6π‘₯ + 2𝑦 + 4𝑧 = 6

a. Apply Gaussian elimination method to reduce the system to triangular form.

b. What do you observe from your answer in part (a) above?


1
Expert's answer
2021-10-04T14:14:10-0400

a.

Augmented matrix is


"\begin{pmatrix}
 3 & 2 & 1 & & 3 \\
 2 & 1 & 1 & & 0 \\
 6 & 2 & 4 & & 6 \\
\end{pmatrix}"

"R_1=R_1\/3"


"\begin{pmatrix}
 1 & 2/3 & 1/3 & & 1 \\
 2 & 1 & 1 & & 0 \\
 6 & 2 & 4 & & 6 \\
\end{pmatrix}"

"R_2=R_2-2R_1"


"\begin{pmatrix}
 1 & 2/3 & 1/3 & & 1 \\
 0& -1/3 & 1/3 & & -2\\
 6 & 2 & 4 & & 6 \\
\end{pmatrix}"

"R_3=R_3-6R_1"


"\begin{pmatrix}
 1 & 2/3 & 1/3 & & 1 \\
 0& -1/3 & 1/3 & & -2\\
 0 & -2 & 2 & & 0 \\
\end{pmatrix}"

"R_2=-3R_2"


"\begin{pmatrix}
 1 & 2/3 & 1/3 & & 1 \\
 0 & 1 & -1 & & 6\\
 0 & -2 & 2 & & 0 \\
\end{pmatrix}"

"R_1=R_1-2R_2\/3"


"\begin{pmatrix}
 1 & 0 & 1 & & -3\\
 0 & 1 & -1 & & 6\\
 0 & -2 & 2 & & 0 \\
\end{pmatrix}"

"R_3=R_3+2R_2"


"\begin{pmatrix}
 1 & 0 & 1 & & -3\\
 0 & 1 & -1 & & 6\\
 0 & 0 & 0 & & 12 \\
\end{pmatrix}"

"R_3=R_3\/12"


"\begin{pmatrix}
 1 & 0 & 1 & & -3\\
 0 & 1 & -1 & & 6\\
 0 & 0 & 0 & & 1 \\
\end{pmatrix}"

"R_1=R_1+3R_3"


"\begin{pmatrix}
 1 & 0 & 1 & & 0\\
 0 & 1 & -1 & & 6\\
 0 & 0 & 0 & & 1 \\
\end{pmatrix}"

"R_2=R_2-6R_3"


"\begin{pmatrix}
 1 & 0 & 1 & & 0\\
 0 & 1 & -1 & & 0\\
 0 & 0 & 0 & & 1 \\
\end{pmatrix}"

b. The system is inconsistent. The given system has no solution.